# Fast Task Planning with Neuro-Symbolic Relaxation

### v1

@@ Abstract
Real-world task planning requires long-horizon reasoning over large sets of entities with complex relationships and attributes, leading to a combinatorial explosion for classical symbolic planners. To prune the search space, recent methods prioritize searching on a simplified task only containing a few “important” entities predicted by a neural network. However, such a simple neuro-symbolic (NeSy) integration risks omitting critical entities and wasting resources on unsolvable simplified tasks. To enable Fast and reliable planning, we introduce a NeSy relaxation strategy (Flax), combining neural importance prediction with symbolic expansion. Specifically, we first learn a graph neural network to predict entity importance to create a simplified task and solve it with a symbolic planner. Then, we solve a rule-relaxed task to obtain a quick rough plan, and reintegrate all referenced entities into the simplified task to recover any overlooked but essential elements. Finally, we apply complementary rules to refine the updated task, keeping it both reliable and compact. Extensive experiments are conducted on both synthetic and real-world maze navigation benchmarks where a robot must traverse through a maze and interact with movable objects. The results show that Flax boosts the average success rate by 20.82% and cuts mean wall-clock planning time by 17.65% compared with the state-of-the-art NeSy baseline. We expect that Flax offers a practical path toward fast, scalable, long-horizon task planning in complex environments.
@@ I Introduction
Classical task planners like Fast Downward [5] offer formal guarantees on problems encoded in Planning Domain Definition Language (PDDL) [6] but scale poorly with the number of entities [^3] Throughout the remainder of this paper we use the term entity in place of PDDL’s standard term object. This change in nomenclature is intentional: in our problem domain, we distinguish three fundamental types, robot, position and object, thus reserving entity to avoid any ambiguity. and relations [7]. For instance, consider a task where a robot must navigate through a maze-like environment while manipulating movable objects, executing primitive actions like move, push, pickup, and place. Each additional operator exponentially increases the branching factor of the action space, inducing a combinatorial explosion in the discrete task planning graph, which can encompass millions of reachable states prior to incorporating continuous motion planning constraints. For example, on a $15\!\times\!15$ maze, a state-of-the-art classical planner can spend several minutes expanding useless nodes even when equipped with advanced heuristics such as $h^{FF}$ [8] and LM-cut [9]. This is an unacceptable delay for many embedded systems.
@@
To address this, two branches of research have emerged. The first learns an efficient neural policy via deep reinforcement learning (DRL) [11]. Despite the promising results, most algorithms remain highly sample-inefficient [12] and generalize poorly on novel states [7]. For PSPACE-hard planning domains [13], the size of policy grows exponentially with the number of entities [14, 15], rendering the policies impractically large for memory-constrained hardware and infeasible for most fixed-size neural networks. The second direction [10, 16, 17] learns a neural predictor that shrinks the search space for symbolic planners.
@@
For instance, PLOI [10] utilizes a Graph Neural Network (GNN) to filter out irrelevant entities, thus accelerating symbolic search. This approach necessitates a symbolic planner to iteratively verify task solvability as the filtering threshold decreases. However, in complex, long-horizon tasks with large search spaces, the planner often requires considerable time to determine unsolvability. Furthermore, due to the absence of correction mechanism, even minor mispredictions can exhaust the entire computational budget. This underscores our key insight: effective handling of complex planning tasks demands a lightweight correction mechanism to promptly detect and correct pruning errors, preventing unnecessary computational overhead.
@@
To this end, we propose Flax, a neuro-symbolic relaxation strategy for Fast task planning that couples neural importance prediction with symbolic rough-plan enhancement and complementary rule expansion. Specifically, we first employ a neural predictor with a short time budget to create an initial simplified task. If the task is not solved, we solve a relaxed version of the full task with domain-specific symbolic relaxation rules; every entity involved by the rough plan is merged back to recover entities that the predictor may have missed. The symbolic search at this step does not aim to solve the full task, but rather to identify potentially important entities, so that pruning errors can be fixed without much effort. Finally, we apply symbolic complementary rules to restore more potential missing entities. This strategy offers a near‑zero‑cost way to recover local consistency after rough-plan enhancement and thus largely preserves overall success rate. Together, they combine the raw speed of neural filtering with the reliability of symbolic reasoning, producing simplified tasks that are far smaller than the original one yet significantly decrease the possibility of dead-end failures.
@@
We evaluate Flax on a new Maze Navigation among movable objects (MazeNamo) benchmark in MiniGrid [18] (three map sizes and four difficulty levels), high-fidelity Isaac Sim [19] forklift-warehouse scenarios, and real-world trials on a Unitree Go2 with a D1 manipulator. Under identical time budgets, Flax boosts the average success rates by 20.82% and cuts mean wall-clock planning time by 17.65% relative to the previous neuro-symbolic baseline, while classical planning alone struggles on most hard and expert tasks, as is shown in Fig. 1. These gains carry over to real-world deployment: the same model drives the quadruped to pick, place, and push objects in cluttered real-world scenes, closing the loop from abstract PDDL reasoning to kinodynamic execution with no additional tuning. We summarize our key contributions as follows:
@@
  • Symbolic Relaxation and Complementation: We propose a domain-specific symbolic relaxation and an entity recovery method to quickly correct neural prediction errors, mitigating planning dead-ends.
@@ II-A Neuro-Symbolic Task Planning for Robots
Task planning produces discrete action sequences that achieve high-level goals in robotics [20], logistics [21], and games [22]. Classical forward-search planners such as Fast Downward [5] offer completeness and optimality guarantees, yet direct applicability to robotics is limited due to (1) They usually assume a fully discrete, factored state space and (2) The search complexity grows exponentially with the number of entities [7]. To overcome these limitations, neuro-symbolic approaches integrate learned neural components into the planning loop. One stream of work leverages neural networks to lift raw states into predicates [23, 24, 25, 26] and raw actions into skills and operators [27, 28, 29]. The resulting hybrid system makes long-horizon task-and-motion planning (TAMP) possible on real robots [27, 24, 30, 26]. A complementary line assumes the planning model is given and fixed, while uses neural networks to make planning more efficient. Early systems learned heuristics or value functions to guide best-first search [31, 22], while more recent methods predict entity-level relevance with graph neural networks (GNNs) [10, 17]. Our work falls within the second stream: assuming a given symbolic domain description and low-level skills, we focus on accelerating planning without modifying the predicates or operators.
@@ II-B Learning for Efficient Task Planning
Learning heuristics and relaxations  A third family learns heuristic estimates or lightweight relaxations that bias classical search. Learned admissible heuristics [31] and value-guided graph search [22] fall into this category. Related works also do graph learning for heuristics in discrete and numeric domains [10, 17]. Our approach also falls into this category: we apply a domain-specific relaxation that yields a rough plan retaining only entities required for optimality, significantly reducing the grounded planning problem size.
@@ III Preliminaries
$\Pi=\langle\mathcal{D},\;\tau,\;T\rangle,$ (1)
where $\mathcal{D}$, $\tau$, and $T$ denote a domain, a task instance, and a transition model, respectively, and will be explained next.
A Domain describes lifted symbols that are shared across all concrete planning tasks. Formally it is defined as:
@@
where $\mathcal{P}$ is a finite set of lifted predicates $p\in\mathcal{P}$, each defined over a list of typed variables; $\mathcal{A}$ is a finite set of lifted actions where each action $\alpha\in\mathcal{A}$ is defined as a quadruple $\alpha=\langle\mathtt{pars}(\alpha),\;\mathtt{pre}(\alpha),\;\mathtt{add}(\alpha),\;\mathtt{del}(\alpha)\rangle$, $\mathtt{pars}(\alpha)$. Here, $\mathtt{pars}(\alpha)$ is a list of typed variables, and $\mathtt{pre}(\alpha)$, $\mathtt{add}(\alpha)$, and $\mathtt{del}(\alpha)$ are sets of lifted predicates, with all free variables drawn from $\mathtt{pars}(\alpha)$. Specifically, if a set of concrete entities is specified for a particular task, grounding the predicates yields the corresponding state space.
Task distribution  $\mathcal{T}$ is defined over tasks $\tau$
@@
where $\mathcal{O}$ is a finite set of typed entities. Following [10], we also treat entity types as unary predicates. We denote $P(\mathcal{O})$ as the set obtained by Grounding all lifted predicates with entities from $\mathcal{O}$. and a state is any subset $s\!\subseteq\mathcal{\!}P(\mathcal{O})$. Thus, $I\!\subseteq\mathcal{\!}P(\mathcal{O})$ is the initial state, and $G\!\subseteq\mathcal{\!}P(\mathcal{O})$ is a set of ground goal atoms that must be made true after plan execution.
Plans  A ground action $a$ is obtained by replacing the variable parameters of a lifted action $\alpha$ with appropriately typed objects from $\mathcal{O}$. We denote by $T$ the deterministic transition function induced by $\mathcal{A}$:
$T(s,a)\;=\;(s\setminus\mathtt{del}(a))\,\cup\,\mathtt{add}(a)\quad\text{iff}\quad\mathtt{pre}(a)\subseteq s.$
A finite sequence of ground actions $\pi=\langle a_{1},a_{2},\ldots,a_{H}\rangle$ is a plan for task $\tau$ iff successive application of $T$ maps $I$ to a terminal state $s_{H}$ with $G\subseteq s_{H}$.
PDDL files  In standard PDDL practice, the domain description $\mathcal{D}$ is stored in a domain file, while a sampled task $\tau$ corresponds to a problem file that lists $\mathcal{O}$, $I$ and $G$.
@@ IV-A MazeNamo Problem Setup
In this work, we define a challenging maze navigation problem among movable objects, where a robot must navigate to a goal position while manipulating both heavy and light boxes. Some general rules and constraints are:
@@
  • Types: ?r - robot, ?o - object, ?p - pos.
  • Key predicates: robot/object position (rAt(?r,?p), oAt(?o,?p)), orientation (dirIsLeft(?r), …), adjacency (upTo(?p,?p), leftTo(?p,?p), …), object category (isHeavy(?o), isLight(?o)), object status (onGround(?o), upon(?o,?o), clear(?o)), and grasp state (handempty(?r), holding(?r,?o)).
  • Action classes: Each abstract capability (Turn, Move, Push, PickUp, and Place) is realized by several concrete operators. For example, Place is split into “on ground” vs. “on object” variants, each expanding to four orientation‑specific actions.
Goal  Each instance specifies a single goal literal rAt(robot, pgoal) (robot reaching a specific position).
Challenges and Illustrative Example  Fig. 2 shows a typical maze. A heavy box (blue) blocks the path to the green goal, bordered by two stacked light boxes (yellow) and walls (grey). The robot must: (i) pick up the front light box, (ii) relocate it, (iii) turn left and push, (iv) lift the second light box, and (v) move into the now-clear goal cell, yielding the 9-step plan in Fig. 2. Compared to Sokoban [38], MazeNamo adds two box categories, stacking, pick-and-place, and orientation constraints, largely expanding the search space and branching factor. Standard heuristic planners thus face harder instances, and neural relaxations often fail on long horizons where search cannot easily recover from mispredictions. On the $15\times 15$ expert-level mazes, the state-of-the-art classical planner Fast Downward takes over 100 seconds on average to find a satisficing plan, which is too slow for real-world applications with strict time budgets.
Figure 3: Flax pipeline. Pre-processing: convert the task to PDDL and encode it as a scene graph $\mathcal{G}$. Training: a symbolic planner provides an optimal plan; entities in the plan are labeled 1 as important, others 0. A GNN learns these labels from $\mathcal{G}$ with BCE loss. Inference (budget $B$): planning runs in three steps. Step 1: keep entities with importance $>q$ to form $\tau_{\mathcal{O}_{1}}$; lower $q$ until a plan $\pi$ is found or time $\Delta t_{1}$ expires. Step 2: if Step 1 fails, solve a rule-relaxed task within $\Delta t_{2}$ to obtain rough plan $\pi_{r}$; merge its entities with $\mathcal{O}_{1}$ to get $\mathcal{O}_{2}$. Step 3: lightweight rules restore consistency, yielding $\mathcal{O}_{3}$. Remaining time plans on $\tau_{\mathcal{O}_{3}}$; the plan is then validated on the original task.
@@ IV-B Problem Formulation
We study the following problem: Within a time budget, maximize the success rate of finding a satisficing plan for a PDDL task guided by a neural importance predictor. This can be formulated as a bilevel optimization [39], where the lower-level optimization trains a neural predictor, and the upper-level maximizes test-time success within the time budget using the neural predictor and a symbolic planner:
$\max_{\bm{\mu}}\;$ $\;\mathcal{U}\bigl{(}s(\bm{\theta}^{*}),\tau;\bm{\mu}\bigr{)},$ (4a)
s. t. $\quad\bm{\theta}^{*}=\arg\min_{\bm{\theta}}\;\mathcal{L}(\tau;\bm{\theta}),$ (4b)
$\quad t\bigl{(}s(\bm{\theta}^{*}),\tau;\bm{\mu}\bigr{)}\leq B,$ (4c)
where $\mathcal{U}$ is the upper-level objective, $s$ is the neural importance predictor with parameters $\bm{\theta}$, $\tau$ is the PDDL task to solve, $\bm{\mu}$ is the parameters of the symbolic modules, $\mathcal{L}$ is the lower-level objective, $t$ is the actual elapsed time used to search for a plan, and $B$ is the total time budget.
Upper-level Objective  Given a learned importance predictor $s({\bm{\theta}^{*}})$, the upper-level objective $\mathcal{U}\in\{0,1\}$ indicates whether a satisficing plan $\pi$ is found within the time budget $B$ using a symbolic planner parameterized by $\bm{\mu}$.
Lower-level Objective  The neural importance predictor $s_{\theta}$ assigns a score $s_{\theta}(o)\in[0,1]$ to every entity node $o$ in the scene graph induced by $\tau$ similar to [10]. Given an optimal plan $\pi^{*}$ for task $\tau$, we assign each node a binary label
$\ell(o)=\begin{cases}1&\text{if }o\text{ appears in }\mathtt{pars}(a)\text{ for any }a\in\pi^{\star},\\[2.0pt] 0&\text{otherwise.}\end{cases}$
@@
$\begin{split}\mathcal{L}\!=\!\frac{-1}{|V|}\sum_{o\in V}\Bigl{[}\ell(o)\log s(o)\!+\!\bigl{(}1\!-\!\ell(o)\bigr{)}\log\bigl{(}1\!-\!s(o)\bigr{)}\Bigr{]},\end{split}$ (5)
where $V$ denotes the entity set of $\tau$.

### v2

@@ Abstract
Real-world task planning requires long-horizon reasoning over large sets of objects with complex relationships and attributes, leading to a combinatorial explosion for classical symbolic planners. To prune the search space, recent methods prioritize searching on a simplified task only containing a few “important” objects predicted by a neural network. However, such a simple neuro-symbolic (NeSy) integration risks omitting critical objects and wasting resources on unsolvable simplified tasks. To enable Fast and reliable planning, we introduce a NeSy relaxation strategy (Flax), combining neural importance prediction with symbolic expansion. Specifically, we first learn a graph neural network to predict object importance to create a simplified task and solve it with a symbolic planner. Then, we solve a rule-relaxed task to obtain a quick rough plan, and reintegrate all referenced objects into the simplified task to recover any overlooked but essential elements. Finally, we apply complementary rules to refine the updated task, keeping it both reliable and compact. Extensive experiments are conducted on both synthetic and real-world maze navigation benchmarks where a robot must traverse through a maze and interact with movable obstacles. The results show that Flax boosts the average success rate by 20.82% and cuts mean wall-clock planning time by 17.65% compared with the state-of-the-art NeSy baseline. We expect that Flax offers a practical path toward fast, scalable, long-horizon task planning in complex environments.
@@ I Introduction
Classical task planners like Fast Downward [13] offer formal guarantees on problems encoded in Planning Domain Definition Language (PDDL) [8] but scale poorly with the number of objects and relations [21]. For instance, consider a task where a robot must navigate through a maze-like environment while manipulating movable obstacles, executing primitive actions like move, push, pickup, and place. Each additional operator exponentially increases the branching factor of the action space, inducing a combinatorial explosion in the discrete task planning graph, which can encompass millions of reachable states prior to incorporating continuous motion planning constraints. For example, on a $15\!\times\!15$ maze, a state-of-the-art classical planner can spend several minutes expanding useless nodes even when equipped with advanced heuristics such as hhFF [14] and LM-cut [12]. This is an unacceptable delay for many embedded systems with strict real-time requirements.
@@
To address this, two branches of research have emerged. The first learns an efficient neural policy via deep reinforcement learning (DRL) [25]. Despite the promising results, most algorithms remain highly sample-inefficient [33] and generalize poorly on novel states [21]. For PSPACE-hard planning domains [2], the size of policy grows exponentially with the number of objects [34, 24], rendering the policies too large for memory-constrained hardware and infeasible for most fixed-size neural networks. The second direction [31, 38, 3] learns a neural predictor that shrinks the search space for symbolic planners.
@@
For instance, PLOI [31] utilizes a Graph Neural Network (GNN) to filter out irrelevant objects, thus accelerating symbolic search. This approach necessitates a symbolic planner to iteratively verify task solvability as the filtering threshold decreases. However, in complex, long-horizon tasks with large search spaces, the planner often requires considerable time to determine unsolvability. Furthermore, due to the absence of correction mechanism, even minor mispredictions can exhaust the entire computational budget. This underscores our key insight: effective handling of complex planning tasks demands a lightweight correction mechanism to promptly detect and correct pruning errors, preventing unnecessary overhead.
@@
To this end, we propose Flax, a neuro-symbolic relaxation strategy for Fast task planning that couples neural importance prediction with symbolic rough-plan enhancement and complementary rule expansion. Specifically, we first employ a neural predictor with a short time budget to create an initial simplified task. If the task is not solved, we solve a relaxed version of the full task with domain-specific symbolic relaxation rules; every object involved by the rough plan is merged back to recover objects that the predictor may have missed. The symbolic search at this step does not aim to solve the full task, but rather to identify potentially important objects, so that pruning errors can be fixed without much effort. Finally, we apply symbolic complementary rules to restore more potential missing objects. This strategy offers a fast local consistency recovery after rough-plan enhancement and thus largely preserves overall success rate. Together, they combine the raw speed of neural filtering with the reliability of symbolic reasoning, producing simplified tasks that are far smaller than the original one yet significantly decrease the possibility of dead-end failures.
@@
We evaluate Flax on a new Maze Navigation among movable obstacles (MazeNamo) benchmark in MiniGrid [4] (three map sizes and four difficulty levels), high-fidelity Isaac Sim [23] forklift-warehouse scenarios, and real-world trials on a Unitree Go2 with a D1 manipulator. Under identical time budgets, Flax boosts the average success rates by 20.82% and cuts mean wall-clock planning time by 17.65% relative to the previous neuro-symbolic baseline, while classical planning alone struggles on most hard and expert tasks, as is shown in Fig. 1. These gains carry over to real-world deployment: the same model drives the quadruped to pick, place, and push obstacles in cluttered real-world scenes, closing the loop from abstract PDDL reasoning to kinodynamic execution with no additional tuning. We summarize our contributions as follows:
@@
  • Symbolic Relaxation and Complementation: We propose a domain-specific symbolic relaxation and an object recovery method to quickly correct neural prediction errors, mitigating planning dead-ends.
@@ II-A Neuro-Symbolic Task Planning for Robots
Task planning produces discrete action sequences that achieve high-level goals in robotics [10], logistics [27], and games [29]. Classical forward-search planners such as Fast Downward [13] offer completeness and optimality guarantees, yet direct applicability to robotics is limited due to (1) They usually assume a fully discrete, factored state space and (2) The search complexity grows exponentially with the number of objects [21]. To overcome these limitations, neuro-symbolic approaches integrate learned neural components into the planning loop. One stream of work leverages neural networks to lift raw states into predicates [20, 32, 11, 22] and raw actions into skills and operators [5, 30, 18]. The resulting hybrid system makes long-horizon task-and-motion planning (TAMP) possible on real robots [5, 32, 17, 22]. A complementary line assumes the planning model is given and fixed, while uses neural networks to make planning more efficient. Early systems learned heuristics or value functions to guide best-first search [1, 29], while more recent methods predict object-level relevance with graph neural networks (GNNs) [31, 3]. Our work falls within the second stream: assuming a given symbolic domain and low-level skills, we focus on accelerating planning without modifying the predicates or operators.
@@ II-B Learning for Efficient Task Planning
Learning heuristics and relaxations  A third family learns heuristic estimates or lightweight relaxations that bias classical search. Learned admissible heuristics [1] and value-guided graph search [29] fall into this category. Related works also do graph learning for heuristics in discrete and numeric domains [31, 3]. Our approach also falls into this category: we apply a domain-specific relaxation that yields a rough plan retaining only objects required for optimality, significantly reducing the grounded planning problem size.
@@ III Preliminaries
$\Pi=\langle\mathcal{D},\;\tau\rangle,$ (1)
where $\mathcal{D}$ and $\tau$ denote a domain and a task instance.
A Domain $\mathcal{D}$ describes lifted symbols that are shared across all concrete planning tasks. Formally it is defined as:
@@
where $\mathcal{P}$ is a finite set of lifted predicates $p\in\mathcal{P}$, each defined over a list of typed variables; $\mathcal{A}$ is a finite set of lifted actions where each action $\alpha\in\mathcal{A}$ is defined as a quadruple $\alpha=\langle\mathtt{pars}(\alpha),\;\mathtt{pre}(\alpha),\;\mathtt{add}(\alpha),\;\mathtt{del}(\alpha)\rangle$, $\mathtt{pars}(\alpha)$. Here, $\mathtt{pars}(\alpha)$ is a list of typed variables, and $\mathtt{pre}(\alpha)$, $\mathtt{add}(\alpha)$, and $\mathtt{del}(\alpha)$ are sets of lifted predicates, with all free variables drawn from $\mathtt{pars}(\alpha)$. Specifically, if a set of concrete objects is specified for a particular task, grounding the predicates yields the corresponding state space.
Task $\tau$ is defined as
@@
where $\mathcal{O}$ is a finite set of typed objects. Following [31], we also treat object types as unary predicates. We denote $P(\mathcal{O})$ as the set obtained by grounding all lifted predicates with objects from $\mathcal{O}$, and a state is any subset $s\!\subseteq\mathcal{\!}P(\mathcal{O})$. Thus, $I\!\subseteq\mathcal{\!}P(\mathcal{O})$ is the initial state, and $G\!\subseteq\mathcal{\!}P(\mathcal{O})$ is a set of grounded goal atoms that must be true after plan execution.
Plans  A grounded action $a$ is obtained by replacing the variable parameters of a lifted action $\alpha$ with appropriately typed objects from $\mathcal{O}$. A finite sequence of grounded actions $\langle a_{1},a_{2},\ldots,a_{H}\rangle$ is a plan for task $\tau$.
@@ IV-A MazeNamo Problem Setup
In this work, we define a challenging maze navigation problem among movable obstacles, where a robot must navigate to a goal position while manipulating both heavy and light boxes. Some general rules and constraints are:
@@
  • Types: ?r - robot, ?o - obstacle, ?p - pos.
  • Key predicates: robot/obstacle position (rAt(?r,?p), oAt(?o,?p)), orientation (dirIsLeft(?r), …), adjacency (upTo(?p,?p), leftTo(?p,?p), …), obstacle category (isHeavy(?o), isLight(?o)), obstacle status (onGround(?o), upon(?o,?o), clear(?o)), and grasp state (handempty(?r), holding(?r,?o)).
  • Action classes: Each abstract capability (Turn, Move, Push, PickUp, and Place) is realized by several concrete operators. For example, Place is split into “on ground” vs. “on obstacles” variants, each expanding to four orientation‑specific actions.
[…]
Challenges and Illustrative Example  Fig. 2 shows a typical maze. A heavy box (blue) blocks the path to the green goal, bordered by two stacked light boxes (yellow) and walls (grey). The robot must: (i) pick up the front light box, (ii) relocate it, (iii) turn left and push, (iv) lift the second light box, and (v) move into the now-clear goal cell, yielding the 9-step plan in Fig. 2. Compared to Sokoban [6], MazeNamo adds two box categories, stacking, pick-and-place, and orientation constraints, largely expanding the search space and branching factor. Standard heuristic planners thus face harder instances, and neural relaxations often fail on long horizons where search cannot easily recover from mispredictions. On the $15\!\times\!15$ expert-level mazes, the classical planner Fast Downward takes over 100 seconds on average to find a satisficing plan, which is too slow for real-world applications with strict time budgets.
Figure 3: Flax pipeline. Pre-processing: convert the task to PDDL and encode it as a relational graph $\mathcal{G}$. Training: a symbolic planner provides an optimal plan; objects in the plan are labeled 1 as important, others 0. A GNN learns these labels from $\mathcal{G}$ with BCE loss. Planning (budget $B$): planning runs in three steps. Step 1: keep objects with importance >q>q to form $\tau_{\mathcal{O}_{1}}$; lower $q$ until a plan $\mu$ is found or time $\Delta t_{1}$ expires. Step 2: if Step 1 fails, solve a rule-relaxed task within $\Delta t_{2}$ to obtain rough plan $\mu_{r}$; merge its objects with $\mathcal{O}_{1}$ to get $\mathcal{O}_{2}$. Step 3: lightweight rules restore consistency, yielding $\mathcal{O}_{3}$. Remaining time plans on $\tau_{\mathcal{O}_{3}}$; the plan is then validated on the original task.
@@ IV-B Problem Formulation
We study the following problem: Training a neural network to predict accurate importance scores for all the objects in a PDDL task. This can be formulated as a bilevel optimization [35], where the lower level minimizes the total action cost using a symbolic planner and produces an optimal plan, and the upper level minimizes the loss between the predicted importance scores from a neural network and the ground truth importance scores derived from the optimal plan:
$\min_{\bm{\theta}}\;$ $\;\mathcal{U}\bigl(s(\bm{\theta},\tau),\bm{\mu}^{*}\bigr),$ (4a)
s. t. $\quad\bm{\mu}^{*}=\arg\min_{\bm{\mu}}\;\mathcal{L}(\bm{\mu};\tau),$ (4b)
where $\mathcal{U}$ is the upper-level objective, $s$ is the neural importance predictor with parameters $\bm{\theta}$, $\tau$ is the PDDL task to solve, $\bm{\mu}$ is the action sequence produced by the symbolic planner, and $\mathcal{L}$ is the lower-level objective.
Upper-level Objective  The neural importance predictor $s_{\theta}$ assigns a score $s_{\theta}(o)\in[0,1]$ to every object node $o$ in the relational graph induced by $\tau$ similar to [31]. Given an optimal plan $\mu^{*}$ for task $\tau$, we assign each node a binary label
$\ell(o)=\begin{cases}1&\text{if }o\text{ appears in }\mathtt{pars}(a)\text{ for any }a\in\mu^{\star},\\[2.0pt] 0&\text{otherwise.}\end{cases}$
@@
$\begin{split}\mathcal{U}\!=\!\frac{-1}{|V|}\sum_{o\in V}\Bigl[\ell(o)\log s(o)\!+\!\bigl(1\!-\!\ell(o)\bigr)\log\bigl(1\!-\!s(o)\bigr)\Bigr],\end{split}$ (5)
where $V$ denotes the object set of $\tau$.
Lower-level Objective  The lower-level problem solves the PDDL task $\tau$ using a symbolic planner, which searches over all feasible action sequences and returns an optimal plan $\mu^{\star}$ under the domain’s plan-cost metric.